# Deformable Beta Splatting

### v1

@@ Abstract.
3D Gaussian Splatting (3DGS) has advanced radiance field reconstruction by enabling real-time rendering. However, its reliance on Gaussian kernels for geometry and low-order Spherical Harmonics (SH) for color encoding limits its ability to capture complex geometries and diverse colors. We introduce Deformable Beta Splatting (DBS), a deformable and compact approach that enhances both geometry and color representation. DBS replaces Gaussian kernels with deformable Beta Kernels, which offer bounded support and adaptive frequency control to capture fine geometric details with higher fidelity while achieving better memory efficiency. In addition, we extended the Beta Kernel to color encoding, which facilitates improved representation of diffuse and specular components, yielding superior results compared to SH-based methods. Furthermore, Unlike prior densification techniques that depend on Gaussian properties, we mathematically prove that adjusting regularized opacity alone ensures distribution-preserved Markov chain Monte Carlo (MCMC), independent of the splatting kernel type. Experimental results demonstrate that DBS achieves state-of-the-art visual quality while utilizing only 45% of the parameters and rendering 1.5x faster than 3DGS-based methods. Notably, for the first time, splatting-based methods outperform state-of-the-art Neural Radiance Fields, highlighting the superior performance and efficiency of DBS for real-time radiance field rendering.
journal: TOG
@@ 1. Introduction
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021; Barron et al., 2021; Müller et al., 2022; Barron et al., 2023) have demonstrated exceptional capabilities in 3D scene representation and novel view synthesis but are hindered by high computational costs and limited applicability in real-time scenarios due to their reliance on neural network-based encoding and dense volumetric ray-marching. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) introduced an explicit representation paradigm that shifts from volumetric ray-marching to point-based splatting. By representing scenes as a set of 3D Gaussian primitives, 3DGS significantly reduces computational overhead, achieving real-time radiance field rendering while maintaining visual quality comparable to earlier NeRF methods. This explicit representation eliminates the need for heavy neural networks and enables direct manipulation of scene geometry and real-time rendering, making 3DGS ideal for applications such as immersive VR exploration (Jiang et al., 2024; Franke et al., 2024; Yu et al., 2024), dynamic scene representation (Yang et al., 2023; Luiten et al., 2023; Wu et al., 2024; Yang et al., 2024b; Guo et al., 2024), 3D content creation (Tang et al., 2024; Yi et al., 2024a; Chung et al., 2023; Yi et al., 2024b), mesh extraction (Guédon and Lepetit, 2023; Huang et al., 2024; Liu et al., 2024; Dai et al., 2024), semantic field distillation (Shi et al., 2023; Zhou et al., 2024; Qin et al., 2024; Qiu et al., 2024; Ye et al., 2024a), and cinematic production (Wang et al., 2024). However, while 3DGS excels in efficiency, its rendering quality still falls short of the latest NeRF advancements (Barron et al., 2023), particularly in capturing fine details and modeling view-dependent colors. To address this gap, a range of extensions has emerged, focusing on improving its visual quality through anti-alising (Yu et al., 2023; Yan et al., 2024; Liang et al., 2024; Song et al., 2024), deblurring (Lee et al., 2024; Peng et al., 2024; Zhao et al., 2024; Seiskari et al., 2024; Oh et al., 2024), appearance modeling (Xu et al., 2024; Svitov et al., 2024; Weiss and Bradley, 2024; Yang et al., 2024a), and revisting densification (Lu et al., 2023; Niemeyer et al., 2024; Zhang et al., 2024; Liu et al., 2024; Bulò et al., 2024; Kheradmand et al., 2024). Despite these efforts, 3DGS-based methods have yet to surpass the state-of-the-art NeRF methods (Barron et al., 2023).
@@
The comprehensive experimental results demonstrate that our approach consistently outperforms both 3DGS-based and NeRF-based methods in visual quality across all benchmarks. Furthermore, our method utilizes only 45% of the parameters required by 3DGS and renders 1.5x faster than Gaussian Splatting-based methods. It is worthy noting that, for the first time, splatting-based methods surpass state-of-the-art NeRF, demonstrating DBS’s superior performance and efficiency in real-time radiance field rendering.
@@ 2.1. Neural Radiance Fields
Beyond addressing aliasing and blurring, splatting kernel modeling has emerged as an attractive direction for improving rendering quality and efficiency. 2DGS introduces a robust 2D Gaussian splatting framework that simplifies computation while producing better geometries via screen-space rendering (Huang et al., 2024). Li et al. proposed 3D Half Gaussian Splatting (3D-HGS), which optimizes the Gaussian representation by restricting it to half-space for better modeling of occlusions and interactions in complex 3D scenes (Li et al., 2024). Chen et al. developed Linear Kernel Splatting, which replaces Gaussian kernels with linear ones, enabling faster computation and maintaining high-fidelity rendering  (Chen et al., 2024). Despite advancements in previous kernel designs, their limited deformability hinders accurate representation of flat surfaces and sharp edges. In contrast, our Beta Kernel offers flexible shape control, adapting to diverse geometries and improving representation fidelity and rendering quality.
@@
Additionally, appearance modeling has been a crucial area for improving the fidelity of 3DGS renderings. Xu et al. introduced Texture-GS, disentangling appearance from geometry by mapping 2D textures onto 3D surfaces via an MLP for high-fidelity editing (Xu et al., 2024). Svitov et al. proposed BBSplat, using textured planar primitives with learnable RGB textures and alpha maps, achieving over 1200 FPS (Svitov et al., 2024). Weiss and Bradley developed Gaussian Billboards, unifying 2D Gaussian Splatting with UV-parameterized textures for expressive scene representations (Weiss and Bradley, 2024). Yang et al. introduced Spec-Gaussian, leveraging an MLP to model anisotropic spherical Gaussians for complex lighting (Yang et al., 2024a). While existing methods focus on either speed (via billboard textures) or fidelity (via MLPs, with slower rendering), we propose Spherical Beta, which explicitly models appearance by decoupling diffuse and specular components, inspired by the Phong Reflection Model (Phong, 1975).
@@
Revisiting optimization and densification methods has proven crucial for improving 3DGS. Techniques such as Scaffold-GS (Lu et al., 2023), RadSplat (Niemeyer et al., 2024), Pixel-GS (Zhang et al., 2024), and AtomGS (Liu et al., 2024) focus on efficient Gaussian placement and density control. Notably, 3DGS-MCMC (Kheradmand et al., 2024) introduces stochastic processes to refine the sampling and placement of Gaussians, enhancing both rendering quality and speed. These strategies ensure effective capture of scene details while balancing quality with computational cost. Building on these advancements, we extended 3DGS-MCMC principles and designed a Kernel-agnostic densification strategy that stabilizes the optimization process while ensuring spatial distribution consistency.
@@ 3. Deformable Beta Splatting
The primary bottleneck in Gaussian Splatting–based approaches lies in their reliance on a fixed, smooth splatting kernel and low-order SH encoding. To overcome this limitation, we first derive a deformable Beta Kernel function. We then illustrate how this flexible kernel can be applied to both geometric splatting and color encoding, enabling the modeling of varying geometric complexities and lighting frequencies, which enhances the overall scene representation. Finally, we address the unique challenges introduced by employing a deformable kernel and detail how our optimization strategy effectively manages these issues.
@@ 3.1. The Beta Kernel
Figure 3. Analogous to the Phong Reflection Model, learnable spherical betas can effectively capture specular highlights with varying sharpness. Smaller $b$ values result in broader reflections, while higher $b$ values correspond to sharper specular highlights.
@@ 3.3. Spherical Beta
3DGS employs Spherical Harmonics (SH) for view-dependent color encoding. For a SH of degree $N$, the feature dimension scales as $3(N+1)^{2}$, resulting in a quadratic increase in parameters. To achieve real-time performance in practical applications, 3DGS is constrained to using low-order SH (e.g.,$N=3$), which only provide smooth view-dependent colors and struggle to model sharp specular highlights effectively. To overcome these issues, we introduce Spherical Beta function, inspired by the Phong Reflection Model (Phong, 1975):
@@
where $A_{m}$ is the ambient term, $D_{m}$ and $(\hat{R}_{m}\cdot\hat{V})^{\alpha_{m}}c_{m}$ models the diffuse and specular components for light source $m$. Here, $\hat{R}_{m}$ is the normalized reflection direction, $\hat{V}$ is the normalized view direction, $\alpha_{m}$ is the shininess coefficient controlling specularity, and $c_{m}$ is the color of light source. We reorganize the equation by separating view-independent and view-dependent parts and combining the view-independent ambient and diffuse terms, yielding:
(8) $c(\hat{V})=c_{0}+\sum_{m\in\mathcal{M}}(\hat{R}_{m}\cdot\hat{V})^{\alpha_{m}}c_{m}.$
Next, we apply the Beta Kernel to express the specular term:
@@
where $c_{0}$ represents the diffuse or base color for the Beta primitive. For each 3D Beta primitive, we introduce $M=|\mathcal{M}|$ light sources. For each light source $m\in\mathcal{M}$, it is paramertized by reflection direction $\hat{R}_{m}$, color $c_{m}$, and shininess $\beta_{m}$. We can observe that its feature dimension scales as $3+6M$, growing linearly with the number of lights and making it more efficient than SH.
@@
Figure 3 illustrates the specular highlights and corresponding reflection blobs for different $b_{m}$ values, demonstrating the capability of Spherical Beta to model both high and low frequency specularities. By integrating bounded support and adaptive shininess, Spherical Beta enhances the ability to model complex lighting interactions and specular effects with fewer parameters, thereby improving both visual fidelity and computational efficiency.
@@ 3.4. Kernel-Agnostic Markov Chain Monte Carlo
Note that we replace the logit noise function from 3DGS-MCMC with Beta Kernel function for a more compact and well-defined noise function. Here, $\lambda_{\epsilon}$ controls the impact of the noise term, ensuring that the primitive is not only optimized through backpropagation of the loss function but also explores alternative solutions for better representation. $\lambda_{\text{lr}}$ represents the learning rate, and $\Sigma_{\eta}$ is the sampled noise. $\mathcal{B}(o_{i};b^{\prime})$ is the Beta Kernel component, which scales the noise based on the opacity of each primitive, with $b^{\prime}=\ln(25)$ to resemble the fall-off behavior of the original logit function.
@@
The error term $\mathcal{O}(o^{2})$ becomes negligible as we employ an opacity regularizer to maintain a small $o$. Therefore, the difference between original distribution and densified distribution diminishes by regularizing opacity regardless of copy times $N$ and splatting kernel $f(x)$. Detailed proof can be found in Figure 7.
@@ 4.1. Experimental Setup
Our implementation builds upon the 3DGS-MCMC framework (Kheradmand et al., 2024), incorporating modifications to the GSplat library (Ye et al., 2024b). Specifically, we integrate the Beta Kernel and Spherical Beta (SB) functions, which are fully explicit and differentiable, implemented in CUDA for optimized performance. By default, we utilize two light sources in Spherical Beta component per primitive (i.e., $sb=2$), unless otherwise specified. To ensure that the Beta Kernel initially resembles the Gaussian Kernel, we set $b$ to zero at initialization. We observe that different scenes exhibit varying convergence speeds when training the model for the default 30,000 iterations. To accommodate these variations, we have integrated a flexible stopping strategy with patience of 10k iterations into the framework, allowing for two training modes: (Ours, 30K) and (Ours, Full). Detailed training times and iteration counts for each scene are provided in Appendix D. For additional hyperparameters and configuration settings, please refer to Appendix C.
@@ 4.2. Results and Comparisons
Beta Splatting demonstrates superior performance for radiance field reconstruction across all metrics over almost all datasets compared to prior methods. Both the 30K and Full training modes consistently outperform previous implicit and explicit state-of-the-art methods. See Appendix D for detailed per-scene analysis. Qualitative visual comparisons in Figure 5 reveal that while Zip-NeRF tends to exhibit floater artifacts, and 3DGS-MCMC struggles with rendering sharp specularities, DBS maintains high fidelity in both smooth and highly detailed regions, effectively capturing intricate lighting interactions and geometric details.
@@
Table 2. Quantitative evaluation of training and rendering efficiency on the Mip-360 dataset
@@
Efficiency Analysis.
To evaluate efficiency, we compare our method with Zip-NeRF (Barron et al., 2023) and 3DGS-MCMC (Kheradmand et al., 2024) in terms of training efficiency, rendering efficiency, and model size on a single NVIDIA RTX 6000 Ada Generation GPU.. The summarized efficiency metrics are presented in Table 2, with additional per-scene details available in Appendix D. Our findings indicate that DBS not only achieves better reconstruction quality but also renders 50% faster than 3DGS-MCMC (Kheradmand et al., 2024) under the same number of primitives. When sb=2, our method uses less than half storage memory compared to 3DGS-MCMC while achieving the state-of-the-art in all metrics. These results highlight effectiveness of DBS in delivering high-quality renderings with enhanced computational efficiency.
Ablation Study.
Table 3 evaluates the impact of various configurations and kernel choices on PSNR and memory usage for the Tanks&Temples and Deep Blending datasets. Transitioning from the Gaussian to Deformable Beta kernel in DBS with MCMC improves PSNR from 24.29 to 24.45 and from 29.67 to 29.88, respectively. Adopting Kernel-Agnostic MCMC further increases PSNR to 24.57 and 30.03. Setting the number of spherical beta (sb) light sources to 2 achieves a balanced PSNR of 24.79 and 30.10 with memory usage of 200.27 MB and 343.32 MB. Additionally, incorporating a flexible stopping strategy slightly enhances PSNR to 24.85 while maintaining 30.12 without additional memory cost. Overall, DBS with Kernel-Agnostic MCMC and sb=2 provides the best trade-off between rendering quality and memory consumption, further improved by the flexible stopping strategy.

### v2

@@ Abstract.
3D Gaussian Splatting (3DGS) has advanced radiance field reconstruction by enabling real-time rendering. However, its reliance on Gaussian kernels for geometry and low-order Spherical Harmonics (SH) for color encoding limits its ability to capture complex geometries and diverse colors. We introduce Deformable Beta Splatting (DBS), a deformable and compact approach that enhances both geometry and color representation. DBS replaces Gaussian kernels with deformable Beta Kernels, which offer bounded support and adaptive frequency control to capture fine geometric details with higher fidelity while achieving better memory efficiency. In addition, we extended the Beta Kernel to color encoding, which facilitates improved representation of diffuse and specular components, yielding superior results compared to SH-based methods. Furthermore, Unlike prior densification techniques that depend on Gaussian properties, we mathematically prove that adjusting regularized opacity alone ensures distribution-preserved Markov chain Monte Carlo (MCMC), independent of the splatting kernel type. Experimental results demonstrate that DBS achieves state-of-the-art visual quality while utilizing only 45% of the parameters and rendering 1.5x faster than 3DGS-MCMC, highlighting the superior performance of DBS for real-time radiance field rendering. Interactive demonstrations and source code are available on our project website: https://rongliu-leo.github.io/beta-splatting/.
Novel View Synthesis, Radiance Fields, 3D Ellipsoidal Betas, Spherical Betas
journal: TOGccs: Computing methodologies Rasterizationccs: Computing methodologies Shape representationsccs: Computing methodologies Reflectance modelingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021; Barron et al., 2021; Müller et al., 2022; Barron et al., 2023) have demonstrated exceptional capabilities in 3D scene representation and novel view synthesis but are hindered by high computational costs and limited applicability in real-time scenarios due to their reliance on neural network-based encoding and dense volumetric ray-marching. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) introduced an explicit representation paradigm that shifts from volumetric ray-marching to point-based splatting. By representing scenes as a set of 3D Gaussian primitives, 3DGS significantly reduces computational overhead, achieving real-time radiance field rendering while maintaining visual quality comparable to earlier NeRF methods. This explicit representation eliminates the need for heavy neural networks and enables direct manipulation of scene geometry and real-time rendering, making 3DGS ideal for applications such as immersive VR exploration (Jiang et al., 2024; Franke et al., 2024; Yu et al., 2024), dynamic scene representation (Yang et al., 2023; Luiten et al., 2023; Wu et al., 2024; Yang et al., 2024b; Guo et al., 2024), 3D content creation (Tang et al., 2024; Yi et al., 2024a; Chung et al., 2023; Yi et al., 2024b), mesh extraction (Guédon and Lepetit, 2023; Huang et al., 2024; Liu et al., 2024; Dai et al., 2024), semantic field distillation (Shi et al., 2023; Zhou et al., 2024; Qin et al., 2024; Qiu et al., 2024; Ye et al., 2024a), and cinematic production (Wang et al., 2024). However, while 3DGS excels in efficiency, its rendering quality still falls short of the latest NeRF advancements (Barron et al., 2023), particularly in capturing fine details and modeling view-dependent colors. To address this gap, a range of extensions has emerged, focusing on improving its visual quality through anti-alising (Yu et al., 2023; Yan et al., 2024; Liang et al., 2024; Song et al., 2024), deblurring (Lee et al., 2024; Peng et al., 2024; Zhao et al., 2024; Seiskari et al., 2024; Oh et al., 2024), appearance modeling (Xu et al., 2024; Svitov et al., 2024; Weiss and Bradley, 2024; Yang et al., 2024a), and revisting densification (Lu et al., 2023; Niemeyer et al., 2024; Zhang et al., 2024; Liu et al., 2024; Bulò et al., 2024; Kheradmand et al., 2024). Despite these efforts, 3DGS-based methods have yet to surpass the state-of-the-art NeRF methods (Kulhanek and Sattler, 2024).
@@
The comprehensive experimental results demonstrate that our approach consistently outperforms both 3DGS-based and NeRF-based methods in visual quality across all benchmarks. Furthermore, our method utilizes only 45% of the parameters required by 3DGS and renders 1.5x faster than 3DGS-MCMC, demonstrating DBS’s superior performance in real-time radiance field rendering.
@@ 2.1. Neural Radiance Fields
Beyond addressing aliasing and blurring, splatting kernel modeling has emerged as an attractive direction for improving rendering quality and efficiency. 2DGS introduced a perspective-accurate 2D splatting pipeline that leverages ray–splat intersections and rasterization (Huang et al., 2024). Li et al. proposed 3D Half Gaussian Splatting, which optimizes the Gaussian representation by restricting it to half-space for better modeling of occlusions and interactions in complex 3D scenes (Li et al., 2024). Hamdi et al. replaced Gaussians with a Generalized Exponential Function, capturing sharp edges with fewer primitives (Hamdi et al., 2024). Chen et al. developed Linear Kernel Splatting, swapping Gaussian kernels for linear ones to accelerate computation while preserving high-fidelity output (Chen et al., 2024). Recent advances in splatting kernels include 3D Convex Splatting (Held et al., 2024), which swaps 3D Gaussians for smooth, convex 3D primitives to better resolve edges and planes, and Deformable Radial Kernel Splatting (Huang et al., 2025), which enriches 2D Gaussians with learnable radial basis functions plus optimized ray–primitive intersection and culling for flexible, boundary-sharp shape modeling. Despite advancements in splatting kernel designs, the absence of an aligned view-dependent color modeling and dedicated densification strategies for deformable kernels constrains further improvements in final rendering quality.
@@
In the absence of a coherent view-dependent color framework, sophisticated appearance modeling becomes essential to elevate 3DGS rendering fidelity. Xu et al. introduced Texture-GS, disentangling appearance from geometry by mapping 2D textures onto 3D surfaces via an MLP for high-fidelity editing (Xu et al., 2024). Svitov et al. proposed BBSplat, using textured planar primitives with learnable RGB textures and alpha maps, achieving over 1200 FPS (Svitov et al., 2024). Weiss and Bradley developed Gaussian Billboards, unifying 2D Gaussian Splatting with UV-parameterized textures for expressive scene representations (Weiss and Bradley, 2024). Yang et al. introduced Spec-Gaussian, leveraging an MLP to model anisotropic spherical Gaussians for complex lighting (Yang et al., 2024a). While existing methods focus on either speed (via billboard textures) or fidelity (via MLPs, with slower rendering), we propose Spherical Beta, which explicitly models appearance by decoupling diffuse and specular components, inspired by the Phong Reflection Model (Phong, 1975).
@@
Likewise, the lack of bespoke kernel densification strategies makes revisiting optimization and densification methods vital for unlocking further performance gains. Techniques such as Scaffold-GS (Lu et al., 2023), RadSplat (Niemeyer et al., 2024), Pixel-GS (Zhang et al., 2024), and AtomGS (Liu et al., 2024) focus on efficient Gaussian placement and density control. Notably, 3DGS-MCMC (Kheradmand et al., 2024) introduces stochastic processes to refine the sampling and placement of Gaussians, enhancing both rendering quality and speed. While these strategies effectively capture scene details and balance rendering quality with computational cost, they are tailored specifically to Gaussian kernels. We therefore extend the principles of 3DGS-MCMC to propose a kernel-agnostic densification strategy that both stabilizes the optimization process and enforces consistent spatial distribution across arbitrary kernel types.
@@ 3. Deformable Beta Splatting
The primary bottleneck in Gaussian Splatting–based approaches lies in their reliance on a fixed, smooth splatting kernel and low-order SH encoding, which inherently hampers their ability to capture flat surfaces and specular highlights. To overcome this limitation, we first derive a deformable Beta Kernel function. We then illustrate how this flexible kernel can be applied to both geometric splatting and color encoding, enabling the modeling of varying geometric complexities and lighting frequencies, which enhances the overall scene representation. Finally, we address the unique challenges introduced by employing a deformable kernel and detail how our optimization strategy effectively manages these issues.
@@ 3.1. The Beta Kernel
Figure 3. Learnable Spherical Beta can effectively capture specular highlights with varying sharpness. Smaller $b$ values result in broader reflections, while higher $b$ values correspond to sharper specular highlights.
@@ 3.3. Spherical Beta
3DGS employs Spherical Harmonics (SH) for view-dependent color encoding. For a SH of degree $N$, the feature dimension scales as $3(N+1)^{2}$, resulting in a quadratic increase in parameters. To achieve real-time performance in practical applications, 3DGS is constrained to using low-order SH (e.g.,$N=3$), which only provide smooth view-dependent colors and struggle to model sharp specular highlights effectively. To address this, we introduce Spherical Beta (SB) function, which is inspired by the Phong Reflection Model (Phong, 1975):
@@
where $A_{m}$ is the ambient term, $D_{m}$ and $(\hat{R}_{m}\cdot\hat{V})^{\alpha_{m}}c_{m}$ models the diffuse and specular components for light source $m$. Here, $\hat{R}_{m}$ is the normalized reflection direction, $\hat{V}$ is the normalized view direction, $\alpha_{m}$ is the shininess coefficient controlling specularity, and $c_{m}$ is the color of light source. Note that the reflection vector $\hat{R}_{m}$ is derived from the surface normal and corresponding light source directions. In contrast, Spherical Beta merges ambient and diffuse into a single base color, then directly models specular lobes via learnable bounded Beta kernels.
@@
where $c_{0}$ represents the diffuse or base color for the Beta primitive. For each 3D Beta primitive, we introduce $M=|\mathcal{M}|$ SB lobes modeling outgoing reflected radiance. For each SB lobe $m\in\mathcal{M}$, it is paramertized by reflection direction $\hat{R}_{m}$, color $c_{m}$, and shininess $b_{m}$. We can observe that its feature dimension scales as $3+6M$, growing linearly with the number of reflection lobes and making it more efficient than SH.
@@
Figure 3 illustrates the specular highlights and corresponding reflection lobes for different $b_{m}$ values, demonstrating the capability of Spherical Beta to model both high and low frequency specularities. By integrating bounded support and adaptive shininess, Spherical Beta enhances the ability to model complex lighting interactions and specular effects with fewer parameters, thereby improving both visual fidelity and computational efficiency.
@@ 3.4. Kernel-Agnostic Markov Chain Monte Carlo
Note that we replace the logit noise function from 3DGS-MCMC with a Beta Kernel function for a more compact and well-defined noise function. Here, $\lambda_{\epsilon}$ controls the impact of the noise term, ensuring that the primitive is not only optimized through backpropagation of the loss function but also explores alternative solutions for better representation. $\lambda_{\text{lr}}$ represents the learning rate, and $\Sigma_{\eta}$ is the sampled noise. $\mathcal{B}(o_{i};b^{\prime})$ is the Beta Kernel component, which scales the noise based on the opacity of each primitive, with $b^{\prime}=\ln(25)$ to resemble the fall-off behavior of the original logit function.
@@
The error term $\mathcal{O}(o^{2})$ becomes negligible as we employ an opacity regularizer to maintain a small $o$. Therefore, the difference between original distribution and densified distribution diminishes by regularizing opacity regardless of copy times $N$ and splatting kernel $f(x)$. Detailed proof can be found in Appendix D.
@@ 4.1. Experimental Setup
Our implementation builds upon the 3DGS-MCMC framework (Kheradmand et al., 2024), incorporating modifications to the GSplat library (Ye et al., 2024b). Specifically, we integrate the Beta Kernel and Spherical Beta (SB) functions, which are fully explicit and differentiable, implemented in CUDA for optimized performance. By default, we utilize two light sources in Spherical Beta component per primitive (i.e., $sb=2$), unless otherwise specified. To ensure that the Beta Kernel initially resembles the Gaussian Kernel, we set $b$ to zero at initialization. We observe that different scenes exhibit varying convergence speeds when training the model for the default 30,000 iterations. To accommodate these variations, we have integrated a flexible stopping strategy with patience of 10k iterations into the framework, allowing for two training modes: (Ours, 30K) and (Ours, Full). Detailed training times and iteration counts for each scene are provided in Appendix B. For additional hyperparameters and configuration settings, please refer to Appendix F.
@@ 4.2. Results and Comparisons
Beta Splatting demonstrates superior performance for radiance field reconstruction across all metrics over almost all datasets compared to prior methods. Both the 30K and Full training modes consistently outperform previous implicit and explicit state-of-the-art methods. See Appendix B for detailed per-scene analysis. Qualitative visual comparisons in Figure 5 reveal that while Zip-NeRF tends to exhibit floater artifacts, and 3DGS-MCMC struggles with rendering sharp specularities, DBS maintains high fidelity in both smooth and highly detailed regions, effectively capturing intricate lighting interactions and geometric details.
Efficiency Analysis.
@@
Table 2. Quantitative evaluation of training and rendering efficiency.
To evaluate efficiency, we compare our method with Zip-NeRF (Barron et al., 2023) and 3DGS-MCMC (Kheradmand et al., 2024) in terms of training efficiency, rendering efficiency, and model size on a single NVIDIA RTX 6000 Ada Generation GPU. The summarized efficiency metrics are presented in Table 2, with additional per-scene details available in Appendix B. Our findings indicate that DBS not only achieves better reconstruction quality but also renders 1.5x faster than 3DGS-MCMC (Kheradmand et al., 2024) under the same number of primitives. When sb=2, our method uses less than half storage memory compared to 3DGS-based methods while achieving the state-of-the-art in all metrics. Additional details on applying compression methods to further reduce storage requirements are provided in Appendix E. These results highlight effectiveness of DBS in delivering high-quality renderings with enhanced computational efficiency.
Ablation Study.
@@
Table 3 compares PSNR and memory usage on the Tanks & Temples and Deep Blending datasets across several DBS configurations: moving from 3DGS-MCMC to a DBS-MCMC raises PSNR from 24.29 dB to 24.45 dB (Tanks & Temples) and from 29.67 dB to 29.88 dB (Deep Blending); switching to Kernel-Agnostic MCMC further boosts PSNR to 24.57 dB and 30.03 dB with no added memory; varying the number of spherical-beta light sources (sb) reveals a trade-off: sb = 1 uses only 155.8 MB but yields 24.53 dB/30.04 dB, sb = 2 hits a sweet spot at 200.3 MB for 24.79 dB/30.10 dB, while sb = 4 or 6 slightly increase PSNR (up to 24.81 dB/30.12 dB) at much higher memory (289–378 MB); finally, adding our flexible stopping strategy to the sb = 2 setup (“DBS (full)”) delivers a final bump to 24.85 dB and 30.12 dB.